Using your dice with two fives and no ones, imagine that you are playing a simple game where you get paid $1 per pip. What is the expected value of a single roll of the die?
Answer #1

P(2) = 1/6 with paid $2

P(3) = 1/6 with $3

P(4) = 1/6 with $4

P(5) = 2/6 with $5

P(6) = 1/6 with $6

(Paid $1 per pip)

Expected value = 2/6 + 3/6 + 4/6 + 10/6 + 6/6

= 25/6

= 4.1666